OpenStudy (anonymous):

@Satellite73 Could you help me with this one as well? 5^(2x+1)=9^(x+1)

OpenStudy (anonymous):

start out the same way what do you get?

OpenStudy (anonymous):

(2x+1)ln5=(x+1)ln9

OpenStudy (anonymous):

multiply out, this time on both sides

OpenStudy (anonymous):

let me know what you get

OpenStudy (anonymous):

and don't be confuse by the \(\ln(5)\) and \(\ln(9)\) treat them like \(a\) and \(b\)

OpenStudy (anonymous):

ln10x+ln5=ln9x+ln9

OpenStudy (anonymous):

oops \(\ln(5)\times 2x=2\ln(5)x\) not \(\ln(10)x\)

OpenStudy (anonymous):

should be \[2\ln(5)x+\ln(5)=\ln(9)x+\ln(9)\] now put each term with an \(x\) in it on one side of the equal side, the ones without an \(x\) on the other it is your choice as to which one to use

OpenStudy (anonymous):

let me know what you get you are two steps away after that one

OpenStudy (anonymous):

2ln5x-ln9x=ln9-ln5

OpenStudy (anonymous):

ok two more steps factor out the \(x\) on the left hand side

OpenStudy (anonymous):

Then factor out an x?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

then divide as the last step

OpenStudy (anonymous):

So x =(ln9-ln5)/(2ln5-ln9)

OpenStudy (anonymous):

yes that should do it. not that hard when you know what to do right?

OpenStudy (anonymous):

I don't completely understand why oops ln(5)×2x=2ln(5)x not ln(10)x

OpenStudy (anonymous):

log is a function, which is why when i write it i always wrote \(\ln(5)\) or \(\ln(x)\) rather than ln5 or lnx

OpenStudy (anonymous):

in other words it is like \(f(x)\) now you would not expect \(f(x)\times 2\) so be equal to \(f(2x)\)

OpenStudy (anonymous):

and it is not true that \(\ln(5)\times 2=\ln(5\times 2)\) then are just different numbers for sure

OpenStudy (anonymous):

you should get in the habit of writing \(\ln(x)\) and when you start trig write \(\sin(x)\) as well if you teacher doesn't do it, it is out of laziness and then is very apt to confuse we even say "log of 5" rather than "log 5"

OpenStudy (anonymous):

That does help me with the confusion. I have been confused since we started this. I really appreciate your help and I WILL start writing it ln(x)

OpenStudy (anonymous):

good, and good luck. once you get the hang of it, it is not that hard also notice that after step 1, which you did on your own, it was algebra from there on in and had nothing really to do with logs